OpenStudy (anonymous):

Leah has two congruent beach balls. The diameter of one ball is 20 inches. Which choice is the closest for the volume of the second ball?

OpenStudy (cwrw238):

volume of sphere = (4/3) * pi* r^3 here r = 20/2 = 10 ins